OpenStudy (anonymous):

7. Simplify the following product of complex numbers. (5 + 3i) · (6 – 2i)

OpenStudy (lalaly):

30-10i+18i +6 = 36+8i

OpenStudy (angela210793):

=30-10i+18i+6=36+8i